Heat of combustion per mole of benzoic acid: -3220kJ/mole

A) The heat capacity of your bomb calorimeter is determined using 1.000g of benzoic acid. When you combust 1.000g in your bomb calorimeter you find the temperature of 1.200kg of water inside the bomb increases from 21.300 degrees C to 23.562 degrees C. What is the heat capacity for this calorimeter in kJ/C?

B) A 1.800 g sample of phenol was burned in your bomb calorimeter. The temp of the calorimeter plus the contents (1.200kg of water plus all the parts of the calorimeter) increased from 21.36 to 26.37 degrees celsius. What is the heat of combustion per gram of phenol?
Answer #1

A)

temperature rise = 23.562 - 21.300 = 2.262 oC

mass of water = 1.200 kg = 1200 g

moles of benzoic acid = 1.000 / 122.12 = 8.19 x 10^-3 mol

Q = 8.19 x 10^-3 x - 3220

   = 26.37 kJ

Q = m Cp dT + Cp dT

26368 = 1200 x 4.184 x 2.262 + Cp x 2.262

Cp = 6636 J / o C

heat capacity = 6.636 kJ/ oC

B)

Q = m Cp dT

   = 1200 x 4.18 x (26.37 - 21.36)

   = 25154 J

heat of combustion = 25.154 / 1.800 = - 13.97 kJ / g

heat of combustion = - 13.97 kJ / g